OpenStudy (anonymous):

Find x and y in terms of a and b for this system. ax+by=0 a^2(x)+b^2(y)=1 also, a does not equal 0, b does not equal 0, and a does not equal b

OpenStudy (anonymous):

Is that second equation:\[a ^{2x} + b ^{y} = 0\]because that is how it is written.

OpenStudy (anonymous):

@Studentc14 ?

OpenStudy (anonymous):

no, sorry, that was confusing the way i posted it. it is a^2(x)+b^2(y)=1

OpenStudy (anonymous):

i fixed it sorry, i didnt see what i had typed. my bad.

OpenStudy (anonymous):

np. Re-writing to find solution then: (a^2)x + (ab)y = 0 adding the negative of equation #1 to equation #2 (a^2)x + (b^2)y = 1 (a^2)x + (ab)y = 0 [(b^2) - (ab)]y = 1 y = 1/[(b^2) - (ab)] Using equation #1, re-written as: x = -by/a x = (-b/a)/[(b^2) - (ab)] = 1/(a^2 - ab)

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

I put these expressions for x and y back into the original equations and they work.

OpenStudy (anonymous):

So that's it. x = 1/(a^2 - ab) and y = 1/(b^2 - ab) Any questions on that?

OpenStudy (anonymous):

(a^2)x + (ab)y = 0 adding the negative of equation #1 to equation #2 (a^2)x + (b^2)y = 1 can you explain this i dont understand what happened

OpenStudy (anonymous):

I made the coefficients on "x" the same so when you subtract equation #1 from equation #2, the "x" term disappears. The right side is just 1 - 0 = 1

OpenStudy (anonymous):

so you multiply the first equation by a?

OpenStudy (anonymous):

yes, because I am using Gaussian elimination and I have to isolate the variables. I just chose "x" so that I came down to one equation in one variable, "y".

OpenStudy (anonymous):

ok. so I was trying to solve this earlier, and i got to here: -by^2+aby+1=0 and then i tried to plug it into the quadratic equation, with -b="a" ab="b" 1="c" would this be a different way to solve it or is this completely off?

OpenStudy (anonymous):

Where did -by^2 + aby + 1 = 0 come from?

OpenStudy (anonymous):

[(b^2) - (ab)]y = 1 you had it like this... i had not factored out a y and my signs were opposite on both sides (i had multiplied by -1 for some reason) so mine was like this -by^2+aby=-1 so then i added the 1 to both sides and it looked like a quadratic

OpenStudy (anonymous):

Before you get too far afield with your explanation, the key to solving this is to recognize that you have 2 equations in 2 variables, so you would be using either elimination or substitution.

OpenStudy (anonymous):

$$ax+by=0\\by=-ax\\a^2x+b^2y=1\\b^2y=1-a^2x\\-abx=1-a^2x\\(a^2-ab)x=1\\x=\frac1{a(a-b)}\\by=-\frac{a}{a(a-b)}=-\frac1{a-b}$$

OpenStudy (anonymous):

Just saw your explanation on where your equation came from. That would be an error because there is no "y^2" term. It all "x" and "y" to the first power.

OpenStudy (anonymous):

... therefore \(x=\frac1{a(a-b)}\), \(y=-\frac1{b(a-b)}\).

OpenStudy (anonymous):

When you are dealing with undefined constants like "a" and "b", it is usually easier to use elimination than substitution. But no quadratic here. Make sense now?

OpenStudy (anonymous):

@oldrin.bataku it would be good to take the negative out of the numerator for y to keep a consistent "1" in the numerator for both "x" and "y".

OpenStudy (anonymous):

Just switch the order of subtraction in the denominator for "y".

OpenStudy (anonymous):

@tcarroll010 it makes no difference, but you could do \(y=\frac1{b(b-a)}\)

OpenStudy (anonymous):

okay i understand this better now. thank you

OpenStudy (anonymous):

True, it can be written either way, but if you keep "1" in the numerator, then you will be able to see a consistent "- ab" term in the denominator for both "x" and "y" and there is a good comparison to be made on how "x" and "y" differ. Anyway, the problem has already been solved.

OpenStudy (anonymous):

Take a look at my 6th post.

OpenStudy (anonymous):

There is a pleasing visual difference of just a^2 for one and b^2 for the other in the denominator.

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @Studentc14

OpenStudy (anonymous):

you're welcome, thank you actually, you were very helpful :)

OpenStudy (anonymous):

Remember, Gaussian elimination is usually the way to go over substitution. And the key was 2 equations in 2 variables. If you get that far in the future, you will generally have clear sailing.

OpenStudy (anonymous):

thanks for the tips!

OpenStudy (anonymous):

The most aesthetic solution IMHO would probably be something like \(x=\frac1ak\), \(y=\frac1bk\) where \(k=\frac1{a-b}\).

OpenStudy (anonymous):

It was a pleasure working with you and I hope I see you again in the problems!